OpenStudy (anonymous):

Determine the length of the major and minor axes of the following ellipses: 41. x^2/16 + y^2/9=1 42. x^2/4 + y^2/49=1 43. x^2+y^2/2=1 44. 4x^2+9y^2=36 45. 8x^2+2y^2=64

OpenStudy (anonymous):

Why do you repost, when the other solver had trying to help and you vanished!

OpenStudy (mertsj):

Oh no!!!

OpenStudy (amistre64):

posting 5 problems in one post seems more like your looking for answers instead of wanting to understand how to get an answer ....

OpenStudy (amistre64):

make sure the right side there is a 1 by dividing it all by whatever that value is then you can read the length of the semi axis from under the x and y parts

OpenStudy (amistre64):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] your axis lengths are then just 2a and 2b

OpenStudy (amistre64):

a 1 hour suspension for abusing the nature of this site. Please take that time to review the code of conduct for appropriate online behaviour, thank you.